LINUX For My Grandma (Annette Maurer)

• Step 1: Get rid of the windowing system: 1. restart computer

2. login: root

3. su
4. cd /home/grandma
5. mv .bash profile .bash profile old
6. mv .bashrc .bashrc old
7. logout

8. login: grandma

9. (You could “throw away” the mouse now; it won’t be

needed in anything below.)

• Step 2: Connect to the Internet:

1. Normally, just type the letter a to connect to the Internet. (We might have to debug this a little, i.e., pick up where we left off when I was in Australia.)

2. When the modem stops buzzing and the messages stop, and you see something like

   June 8 01:22:28 form pppd[15690]: remote IP address 192.168.254.1

you are connected. Press the Ctrl and c keys together to get back the familiar shell prompt

       [grandma@grandma grandma]$

• Step 3: Connect to my computer at Harvard:

1. Type

     ssh modular.fas.harvard.edu

or

     ssh 140.247.28.132

then type your password, which is roxy3584.

2. To make this a single command (modular), type

      echo "ssh 140.247.28.132" > modular
      chmod +x modular

In future, you can just type modular to connect to my computer.

3. If you are correctly connect to the internet, you will be asked for your password. Enter roxy3584.

    grandma@???'s password:  roxy3584

If no password query appears, you are not connected to the internet. Press Ctrl and c together, then go to Step 2.

• Step 4: Read email.

1. On my computer, modular, type pine:

    [grandma@modular grandma]$ pine

2. The program pine is a program for sending and receiving email. Maybe you can figure out how to use it? If not, try typing ? for help. You won’t need the mouse for anything. Type L to see a list of folders, and select the INBOX folder to read your mail. While reading a message type r to reply, follow the prompts, type the message, then type Ctrl and x at the same time to send the message. (The commands are listed on the bottom, and ^X means “hold Ctrl and x at the same time.) This is all you need to know!

3. If you are this far, then we are “home free”, and I’ll teach

you more later.

• Step 5: Shut down the computer properly.

1. Press the Ctrl, Alt, and Delete keys all at the same time.

Wait.

• Some commands we might need while talking (all in lower case)

1. cd
2. chmod +x
3. cp
4. echo

5. ln (“ell enn”)

6. logout

7. ls (“ell ess”)

8. more
9. pine

10. rm

11. ssh

12. su

13. tail

14. >